OpenStudy (ashleyisakitty):

A salsa recipe uses pepper, onion, and tomato in the extended ratio of 3:7:11. How many cups of peppers are needed to make 189 cups of salsa?

OpenStudy (nincompoop):

multiply the cup of salsa by the ratio

OpenStudy (nincompoop):

@abb0t

OpenStudy (nincompoop):

the thing is that I don't know how much (how many cups) a three cup of salsa 3 cups of tomato, 7 cups of onion (that's a lot) and 11 cups of tomato make

OpenStudy (nincompoop):

do you understand @Ashleyisakitty

OpenStudy (ashleyisakitty):

"multiply the cup of salsa by the ratio" do you mean like 189 by 3:7:11? No i dont really understand what you mean.

OpenStudy (nincompoop):

the question is a little to crazy for me to comprehend. if we were to assume that the ratio yields a cup of salsa, then you multiply the cup of salsa to be made by 3

OpenStudy (nincompoop):

189 * 3

OpenStudy (ashleyisakitty):

Hm.. its a multiple choice question and that doesnt fit, i tried that originally.

OpenStudy (nincompoop):

but if the ratio (3:7:11) yields 11 cups of salsa then it's different

OpenStudy (nincompoop):

see what we are missing is the information of how many cups of salsa the ratio 3:7:11 yields

OpenStudy (nincompoop):

realistically, it can't be 1 cup

OpenStudy (abb0t):

How much does the ratio 3:7:11 yield? approximately 1 cup? I think that's what he's asking.

OpenStudy (nincompoop):

I don't think so. it is asking how many cups of pepper would it need to make 189 cups of salsa

OpenStudy (ashleyisakitty):

The question doesnt include how much 3:7:11 could yield, so id assume on the lack of information it would 3:7:11 yield just a cup.

OpenStudy (abb0t):

If it equals 1 cup, then I would just multiply to get an appximate value.

OpenStudy (nincompoop):

but that will lead you to the same answer of 189 * 3 which is incorrect you said I am assuming that the ratio yields you 11 cups if that were the case then you divide 189 by 11 then multiply that by the pepper ratio

OpenStudy (nincompoop):

3 cups of pepper yields you 11 cups of salsa, then 189 cups of salsa requires you ___ cups of pepper

OpenStudy (ashleyisakitty):

51.3 isnt a proper answer choice :/

OpenStudy (nincompoop):

what are the choices?

OpenStudy (ashleyisakitty):

I can choice between 9, 27, 63 and 99.

OpenStudy (ashleyisakitty):

im going to call my teacher lol

OpenStudy (nincompoop):

ya do that LOL or I am just being dumb at the moment and can't see what is being asked and what are the given

OpenStudy (mathstudent55):

Add 3 + 7 + 11 and extend the ratio to one more number. 3 + 7 + 11 = 21 The new even more extended ratio is: P : O : Tom : Total 3 : 7 : 11 : 21 x : a : b : 189 \(\dfrac{3}{x} = \dfrac{21}{189} \)

OpenStudy (ashleyisakitty):

Ah, awesome! So the correct amount of pepper is nine?

OpenStudy (mathstudent55):

No. Nine is the multiplication nfactor, not the final answer. Notice that from the ratio, and using cups, the amount of salsa made is 21 cups. If you make 189 cups, you are making 189/21 = 9 times the amount the ratio states. That means you need to use 9 times the amount of each ingredient. Since the ratio has 3 cups of peppers, you will need 3 * 9 = 27 cups of peppers.

OpenStudy (mathstudent55):

Solving the ratio above will also give you the answer: \( \dfrac{3}{x} = \dfrac{21}{189} \) Cross multiply: 21x = 3 * 189 21x = 567 x = 27

OpenStudy (ashleyisakitty):

Okay, i got it now. Thank you :~)

OpenStudy (mathstudent55):

You're welcome.